# Bimanual Regrasp Planning and Control for Eliminating Object Pose Uncertainty

### v1

@@ V Pose Estimation Using Conformed Grasp Poses
Next, we present the method for estimating the object’s pose. In the planning simulation, we used an assumed object pose. This pose serves as the basis for all grasping and planning processes. Since there is no uncertainty in the simulation, the final object pose remains deterministic. Here, we denote the object’s final pose in the planning simulation as $\boldsymbol{p}_{o_{s}}$ and $\boldsymbol{R}_{o_{s}}$, and denote the three consecutive grasping poses of the robot in the simulation as $\boldsymbol{g}_{s_{1}}$, $\boldsymbol{g}_{s_{2}}$, and $\boldsymbol{g}_{s_{3}}$. Each $\boldsymbol{g}_{s_{i}}$ consists of ($\boldsymbol{p}_{s_{i}}$, $\boldsymbol{R}_{s_{i}}$, $d_{s_{i}}$), where $\boldsymbol{p}_{s_{i}}$ and $\boldsymbol{R}_{s_{i}}$ represent the position and rotation of the grasping center point, and $d_{s_{i}}$ denotes the jaw opening width. On the other hand, by executing the planned motion in the real world, we obtain three consecutive grasping poses of the bimanual robot in the real world as $\boldsymbol{g}_{r_{1}}$, $\boldsymbol{g}_{r_{2}}$, and $\boldsymbol{g}_{r_{3}}$. Each $\boldsymbol{g}_{r_{i}}$ similarly consists of ($\boldsymbol{p}_{r_{i}}$, $\boldsymbol{R}_{r_{i}}$, $d_{r_{i}}$). The first grasping pose $\boldsymbol{g}_{r_{1}}$ aligns exactly with the planned result, while the second grasping pose $\boldsymbol{g}_{r_{2}}$ and the third grasping pose, $\boldsymbol{g}_{r_{3}}$, are adjusted through admittance control to conform to the object’s actual pose. They are different from $\boldsymbol{g}_{s_{2}}$ and $\boldsymbol{g}_{s_{3}}$. Our objective is to estimate the real-world object pose $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ based on these known values.
To address this issue, let us first consider the simulation scenario. Since our three grasping poses involve planar contacts and are mutually orthogonal, the object constrained by them should have a unique pose. The pose can be computed as follows. For each grasp, we can obtain a corresponding opening and closing vector $\boldsymbol{v}_{i}=\boldsymbol{R}_{s_{i}}(\boldsymbol{y})$. Here, $\boldsymbol{R}_{s_{i}}(\boldsymbol{y})$ denotes the $\boldsymbol{y}$ axis (2nd column) of the rotation matrix. We assume the robot hand’s opening and closing direction is aligned with its local $\boldsymbol{y}$ axis. Since the three grasping poses are mutually orthogonal, they uniquely determine the following rotation matrix
$\boldsymbol{R}_{g_{s}}=\begin{bmatrix}\boldsymbol{R}_{s_{1}}(\boldsymbol{y})&\boldsymbol{R}_{s_{2}}(\boldsymbol{y})&\boldsymbol{R}_{s_{3}}(\boldsymbol{y})\end{bmatrix}.$ (3)
Meanwhile, for each grasp, its two parallel fingers define a midplane. Given that the three grasps are mutually orthogonal, their corresponding midplanes intersect at a unique intersection point. This intersection represents a unique position constraint imposed by the grasping configuration. It can be computed using
$\boldsymbol{p}_{g_{s}}=(\boldsymbol{R}_{g_{s}}^{\top})^{-1}\begin{bmatrix}\boldsymbol{R}_{s_{1}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{1}}}\\ \boldsymbol{R}_{s_{2}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{2}}}\\ \boldsymbol{R}_{s_{3}}(\boldsymbol{y})\cdot\boldsymbol{p}_{{s_{3}}}\end{bmatrix}$ (4)
The ($\boldsymbol{p}_{g_{s}}$, $\boldsymbol{R}_{g_{s}}$) obtained from the above two equations are the unique pose determined by the three grasping poses in the planning simulation. Note that it is not the object’s final pose ($\boldsymbol{p}_{o_{s}}$, $\boldsymbol{R}_{o_{s}}$). However, since both are uniquely determined. There is a unique offset between them written as
$\boldsymbol{p}_{\text{diff}}=\boldsymbol{p}_{o_{s}}-\boldsymbol{p}_{g_{s}},~{}\boldsymbol{R}_{\text{diff}}=\boldsymbol{R}_{g_{s}}^{\top}\boldsymbol{R}_{o_{s}}$ (5)
Fig. 6 illustrates the three planes formed by three grasping poses and their intersection point.
Figure 6: Three orthogonal grasping poses with planar contacts determine three planes that intersect at a unique point. Thus, a unique position $\boldsymbol{p}_{g_{s}}$ is determined. Meanwhile, the opening-closing directions of the three gasping poses help constrain a unique rotation $\boldsymbol{R}_{g_{s}}$
In the real world, $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ are uniquely constrained by the grasping poses and have the same offset from the pose determined by the grasping poses. That is
$\boldsymbol{p}_{o_{r}}=\boldsymbol{p}_{g_{r}}+\boldsymbol{p}_{\text{diff}},~{}\boldsymbol{R}_{o_{r}}=\boldsymbol{R}_{g_{r}}\boldsymbol{R}_{\text{diff}}.$ (6)
Thus, $\boldsymbol{p}_{o_{r}}$ and $\boldsymbol{R}_{o_{r}}$ can be computed. Here, $\boldsymbol{p}_{g_{r}}$ and $\boldsymbol{R}_{g_{r}}$ can be obtained similarly to equations (4) and (3) using ($\boldsymbol{p}_{r_{1}}$, $\boldsymbol{R}_{r_{1}}$), ($\boldsymbol{p}_{r_{2}}$, $\boldsymbol{R}_{r_{2}}$), and ($\boldsymbol{p}_{r_{3}}$, $\boldsymbol{R}_{r_{3}}$).
@@ VI Experiments and Analysis
This section presents the real-world validation of the proposed method. Our experimental platform consists of two UR3e robots from Universal Robots. Their installation setup is shown in Fig. 7. The UR3e series robots are equipped with high-precision force/torque sensors at their end effectors, which facilitate the implementation of admittance control. Each robot is installed with a Robotiq Hand-E gripper from Robotiq. The opening range of the gripper is 0$\sim$50 mm. The regrasp planning is performed using the WRS simulation environment Note 2 Note 2 Note 2https://github.com/wanweiwei07/wrs. Fig. 6 in the previous section provided a view of the simulation environment. The coordinate system of our experimental platform is defined with its origin at the intersection of the central plane between the two robot arms and the rear edge of the table, as illustrated by the colored frame in Fig. 7. The $\boldsymbol{x}$-axis points forward, the $\boldsymbol{y}$-axis points towards the left arm, and the $\boldsymbol{z}$-axis aligns with the normal direction of the tabletop surface.

### v3

@@ V Pose Estimation Using Conformed Grasp Poses
V-A Error Parameterization
We denote the pose of the object and first grasp in the world frame of the simulation as $({}^{w}\mathbf{p}_{o}^{\text{sim}},{}^{w}\mathbf{R}_{o}^{\text{sim}})$, and $({}^{w}\mathbf{p}_{g_{1}}^{\text{sim}},{}^{w}\mathbf{R}_{g_{1}}^{\text{sim}})$, respectively. The relative pose of the object with respect to the first grasp can be computed as
{${}^{g_{1}}\mathbf{R}_{o}^{\text{sim}}&=({}^{w}\mathbf{R}_{g_{1}}^{\text{sim}})^{\top}\,{}^{w}\mathbf{R}_{o}^{\text{sim}}\\ {}^{g_{1}}\mathbf{p}_{o}^{\text{sim}}&=({}^{w}\mathbf{R}_{g_{1}}^{\text{sim}})^{\top}({}^{w}\mathbf{p}_{o}^{\text{sim}}-{}^{w}\mathbf{p}_{g_{1}}^{\text{sim}})\end{aligned}\right..$ (3)
Meanwhile, we denote the first grasp’s actual pose after impedance control as $({}^{w}\mathbf{p}_{g_{1}}^{\text{real}},{}^{w}\mathbf{R}_{g_{1}}^{\text{real}})$. Without considering uncertainty, the pose of the object in the world frame of the real world can be computed as
{${}^{w}\mathbf{R}_{o}^{\text{ideal}}({g_{1}})&={}^{w}\mathbf{R}_{g_{1}}^{\text{real}}\cdot{}^{g_{1}}\mathbf{R}_{o}^{\text{sim}}\\ {}^{w}\mathbf{p}_{o}^{\text{ideal}}({g_{1}})&={}^{w}\mathbf{p}_{g_{1}}^{\text{real}}+{}^{w}\mathbf{R}_{g_{1}}^{\text{real}}\cdot{}^{g_{1}}\mathbf{p}_{o}^{\text{sim}}\end{aligned}\right.,$ (4)
The subscript $(g_{1})$ indicates the computations are based on the first grasp. The superscript “ideal” indicates the results are an estimation without considering uncertainty.
To account for uncertainty introduced during the grasping process, we model the object pose error as comprising a rotational component $\mathbf{R}_{o}^{\text{err}}({g_{1}})$ and a translational component $\mathbf{p}_{o}^{\text{err}}({g_{1}})$. Similarly, $(g_{1})$ indicates that the errors are associated with the first grasp. Due to the geometric constraints between the gripper and the object surface, $\mathbf{R}_{o}^{\text{err}}({g_{1}})$ is constrained to a rotation about the second column of the real-world grasp frame, ${}^{w}\mathbf{r}_{g_{1,2}}^{\text{real}}={}^{w}\mathbf{R}_{g_{1}}^{\text{real}}[:,1]$, and $\mathbf{p}_{o}^{\text{err}}({g_{1}})$ is restricted to the plane spanned by the first and third columns ${}^{w}\mathbf{r}_{g_{1,1}}^{\text{real}}={}^{w}\mathbf{R}_{g_{1}}^{\text{real}}[:,0]$ and ${}^{w}\mathbf{r}_{g_{1,3}}^{\text{real}}={}^{w}\mathbf{R}_{g_{1}}^{\text{real}}[:,2]$. The three arrows in Fig. 6(a.ii) show these columns. The errors can thus be parameterized as
$\mathbf{R}_{o}^{\text{err}}({g_{1}})&=\exp\left(\theta\,[{}^{w}\mathbf{r}_{g_{1,2}}^{\text{real}}]_{\times}\right)\\ \mathbf{p}_{o}^{\text{err}}({g_{1}})&=\delta_{g_{1,1}}\,{}^{w}\mathbf{r}_{g_{1,1}}^{\text{real}}+\delta_{g_{1,3}}\,{}^{w}\mathbf{r}_{g_{1,3}}^{\text{real}}\end{aligned}\right..$ (5)
Accordingly, the true pose of the object in the real world is
{${}^{w}\mathbf{R}_{o}^{\text{real}}&={}^{w}\mathbf{R}_{o}^{\text{ideal}}(g_{1})\cdot\mathbf{R}_{o}^{\text{err}}({g_{1}})\\ {}^{w}\mathbf{p}_{o}^{\text{real}}&={}^{w}\mathbf{p}_{o}^{\text{ideal}}(g_{1})+\mathbf{p}_{o}^{\text{err}}({g_{1}})\end{aligned}\right..$ (6)
[…]
V-B Rotation Error Identification
We can use the second grasp to identify $\mathbf{R}_{o}^{\text{err}}({g_{1}})$. Let $({}^{w}\mathbf{p}_{g_{2}}^{\text{sim}},{}^{w}\mathbf{R}_{g_{2}}^{\text{sim}})$ denote the simulated pose of the second grasp (equally the grasp pose before admittance control), and let $({}^{w}\mathbf{p}_{g_{2}}^{\text{real}},{}^{w}\mathbf{R}_{g_{2}}^{\text{real}})$ denote its real-world pose after admittance control. The rotation difference from the simulation to the real execution is $\mathbf{R}_{g_{2}}^{\text{diff}}=({}^{w}\mathbf{R}_{g_{2}}^{\text{sim}})^{\top}\cdot{}^{w}\mathbf{R}_{g_{2}}^{\text{real}}$, which captures the rotational deviation introduced by the admittance control. Since $\mathbf{R}_{o}^{\text{err}}({g_{1}})$ is constrained to only occur around $\mathbf{r}_{g_{1,2}}^{\text{real}}$, we extract the component of $\mathbf{R}_{g_{2}}^{\text{diff}}$ along $\mathbf{r}_{g_{1,2}}^{\text{real}}$ to get $\theta$:
$\theta=\text{axis\_angle}(\mathbf{R}_{g_{2}}^{\text{err}},{}^{w}\mathbf{r}_{g_{1,2}}^{\text{real}}),$ (7)
and consequently $\mathbf{R}_{o}^{\text{err}}({g_{1}})$. The object’s true rotation in the real world can be uniquely identified as
${}^{w}\mathbf{R}_{o}^{\text{real}}={}^{w}\mathbf{R}_{o}^{\text{ideal}}({g_{1}})\cdot\mathbf{R}_{o}^{\text{err}}({g_{1}})={}^{w}\mathbf{R}_{o}^{\text{ideal}}({g_{1}})\cdot\exp\left(\theta\,[{}^{w}\mathbf{r}_{g_{1,2}}^{\text{real}}]_{\times}\right).$ (8)
The rotation of the purple object in Fig. 6(a.iii) illustrates an identified ${}^{w}\mathbf{R}_{o}^{\text{real}}$.
V-C Translation Error Identification
Next, we address $\mathbf{p}_{o}^{\text{err}}({g_{1}})$. In simulation, the pose of the object relative to the second grasp is
{${}^{g_{2}}\mathbf{R}_{o}^{\text{sim}}&=({}^{w}\mathbf{R}_{g_{2}}^{\text{sim}})^{\top}\,{}^{w}\mathbf{R}_{o}^{\text{sim}}\\ {}^{g_{2}}\mathbf{p}_{o}^{\text{sim}}&=({}^{w}\mathbf{R}_{g_{2}}^{\text{sim}})^{\top}({}^{w}\mathbf{p}_{o}^{\text{sim}}-{}^{w}\mathbf{p}_{g_{2}}^{\text{sim}})\end{aligned}\right..$ (9)
Let $({}^{w}\mathbf{p}_{g_{2}}^{\text{real}},{}^{w}\mathbf{R}_{g_{2}}^{\text{real}})$ be the second grasp’s real-world pose. If there is no translational uncertainty, the position of the object in the real world can be computed using the second grasp as
${}^{w}\mathbf{p}_{o}^{\text{ideal}}({g_{2}})={}^{w}\mathbf{p}_{g_{2}}^{\text{real}}+{}^{w}\mathbf{R}_{g_{2}}^{\text{real}}\,{}^{g_{2}}\mathbf{p}_{o}^{\prime}.$ (10)
Since ${}^{w}\mathbf{R}_{o}^{\text{real}}$ has already been identified, the corrected relative position ${}^{g_{2}}\mathbf{p}_{o}^{\prime}$ instead of ${}^{g_{2}}\mathbf{p}_{o}^{\text{sim}}$ is used in the equation. ${}^{g_{2}}\mathbf{p}_{o}^{\prime}$ is computed as
${}^{g_{2}}\mathbf{p}_{o}^{\prime}={}^{g_{2}}\mathbf{R}_{o}^{\text{real}}\left({}^{g_{2}}\mathbf{R}_{o}^{\text{sim}}\right)^{\top}{}^{g_{2}}\mathbf{p}_{o}^{\text{sim}},$ (11)
where
${}^{g_{2}}\mathbf{R}_{o}^{\text{real}}=({}^{w}\mathbf{R}_{g_{2}}^{\text{real}})^{\top}\,{}^{w}\mathbf{R}_{o}^{\text{real}}.$ (12)
The translational uncertainty caused by the second grasp lies within the plane spanned by the first and third columns of ${}^{w}\mathbf{R}_{g_{2}}^{\text{real}}$ and can be parameterized as
$\mathbf{p}^{\text{err}}({g_{2}})=\delta_{2,1}\,{}^{w}\mathbf{r}_{g_{2,1}}^{\text{real}}+\delta_{2,3}\,{}^{w}\mathbf{r}_{g_{2,3}}^{\text{real}},$ (13)
where ${}^{w}\mathbf{r}_{g_{2,1}}^{\text{real}}={}^{w}\mathbf{R}_{g_{2}}^{\text{real}}[:,0]$ and ${}^{w}\mathbf{r}_{g_{2,3}}^{\text{real}}={}^{w}\mathbf{R}_{g_{2}}^{\text{real}}[:,2]$. Fig. 6(a.iii) illustrates them. The true position of the object can be expressed using $\mathbf{p}^{\text{err}}({g_{2}})$ as
${}^{w}\mathbf{p}_{o}^{\text{real}}={}^{w}\mathbf{p}_{o}^{\text{ideal}}(g_{2})+\mathbf{p}_{o}^{\text{err}}({g_{2}}).$ (14)
It is important to note that the translational errors associated with the first and second grasps are respectively constrained to planes determined by $({}^{w}\mathbf{r}_{g_{1,1}}^{\text{real}},{}^{w}\mathbf{r}_{g_{1,3}}^{\text{real}})$ and $({}^{w}\mathbf{r}_{g_{2,1}}^{\text{real}},{}^{w}\mathbf{r}_{g_{2,3}}^{\text{real}})$. Their normal vectors are $\mathbf{n}_{g_{1}}={}^{w}\mathbf{r}_{g_{1,1}}^{\text{real}}\times{}^{w}\mathbf{r}_{g_{1,3}}^{\text{real}}$ and $\mathbf{n}_{g_{2}}={}^{w}\mathbf{r}_{g_{2,1}}^{\text{real}}\times{}^{w}\mathbf{r}_{g_{2,3}}^{\text{real}}$. Since the translational error must simultaneously satisfy the constraints imposed by both planes, it is restricted to lie along their intersection line. The direction of this line can be computed using
$\mathbf{d}_{\text{allowed}}=\frac{\mathbf{n}_{g_{1}}\times\mathbf{n}_{g_{2}}}{\|\mathbf{n}_{g_{1}}\times\mathbf{n}_{g_{2}}\|}.$ (15)
Thus, the translational error can essentially be parameterized using a single scalar value as $\epsilon\,\mathbf{d}_{\text{allowed}}$, and the true real-world object position simplifies to
${}^{w}\mathbf{p}_{o}^{\text{real}}={}^{w}\mathbf{p}_{o}^{\text{ideal}}({g_{2}})+\epsilon\,\mathbf{d}_{\text{allowed}}.$ (16)
The scalar value $\epsilon$ can be easily determined from the displacement of the third grasp after admittance control. Let ${}^{w}\mathbf{p}_{g_{3}}^{\text{sim}}$ and ${}^{w}\mathbf{p}_{g_{3}}^{\text{real}}$ denote the simulated position of the third grasp (position before admittance control) and the real-world position after admittance control, respectively. The displacement is computed as ${}^{w}\mathbf{p}_{g_{3}}^{\text{real}}-{}^{w}\mathbf{p}_{g_{3}}^{\text{sim}}$. The scalar $\epsilon$ can be obtained using the following projection
$\epsilon=\mathbf{d}_{\text{allowed}}^{\top}({}^{w}\mathbf{p}_{g_{3}}^{\text{real}}-{}^{w}\mathbf{p}_{g_{3}}^{\text{sim}})),$ (17)
as indicated within the circle in Fig. 6(a.iv). Substituting the obtained $\epsilon$ into equation (16) yields the object’s true position.
In addition to the L-shaped object with a square cross-section discussed thus far, Fig. 6(b) presents an L-shaped object with a diamond cross-section. Although this object does not contain any pair of strictly orthogonal grasping directions, it is still possible to select three non-parallel grasps and estimate the final object pose using the method described in this section. Fig. 6(b.i) shows the uncertainty after the first grasp, while Fig. 6(b.ii) illustrates the orientation determined after the second grasp, along with the direction of the remaining translational error. A detailed 3D visualization of both the L-shaped objects can be found in the supplementary video.
@@ VI Experiments and Analysis
Our experimental platform consists of two UR3e robots. Their setup is shown in Fig. 7. Each robot is installed with a Robotiq Hand-E gripper. The opening range of the gripper is 0$\sim$50 mm. The regrasp planning is performed using the WRS simulation environment. The coordinate system of the platform is defined at the back of the robot, with its origin at the intersection of the central plane between the two robot arms and the ground. The $\boldsymbol{x}$-axis points toward the monitor in the figure. The $\boldsymbol{y}$-axis points towards the first arm. The $\boldsymbol{z}$-axis points upward. For comparison, we attached optical markers to both the objects and the robots, and tracked the poses of the resulting rigid bodies using three tracking cameras. The poses estimated by the proposed method were compared with those captured by the optical tracking system to evaluate the method’s performance. The tracking system provides a detection accuracy of $\pm$0.5 mm.